Fermi National Laboratory


Accelerator Update for May 9, 2005 - May 11, 2005

Monday May 9
The day shift began with Operations monitoring store 4134 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and with the MiniBooNE, NuMI, and Meson MCenter experiments taking beam.

At 10:24 AM, a vacuum burst in the Main Injector (MI) closed a number of beam valves. It took thirteen minutes to reset everything and have vacuum return to normal.

Operations turned off MiniBooNE beam at 11:04 AM in preparation for an access. They came out of access at 3:54 PM.

At 4:53 PM, Operations reported losing the MI beam permit due to a MECAR trip. This happened two more times over the next seven minutes.

A Tevatron (TeV) expert began an end-of-store study at 5:04 PM.

Operations terminated store 4134 at 5:19 PM.

Operations put the TeV into shot setup at 5:49 PM.

Operations established combination store 4135 at 8:18 PM with an initial luminosity of 84.04E30. Pbar resumed stacking and MCenter resumed taking beam.

A NuMI expert began aperture scans at 8:20 PM.

At 9:11 PM, the Booster East Anode Power Supply (EAPS) tripped off. Operators investigated and had the supply back on fifteen minutes later.

At 11:06 PM, EAPS tripped off again. Operators had it back on ten minutes later.

Tuesday May 10
The midnight shift began with Operations monitoring store 4135, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

At 12:23 AM, EAPS tripped off. Operators had tried to turn on a Booster RF station (BRF19), which caused the trip. They reset the supply, but then it tripped off again at 12:48 AM. Operators reset the supply again.

At 12:55 AM, the NuMI beam permit tripped off. Operators weren't sure of the cause due to a readback problem. Operators called in experts who rebooted a computer. The readbacks showed a MINOS enclosure interlock that had dropped due to a bad module, which they replaced at 5:38 AM.

Operations put the Recycler into shot setup to stash antiprotons at 3:15 AM.

At 3:16 AM, EAPS tripped off for the first of two times in ten minutes. It reset okay both times.

The Recycler stash ended and Pbar resumed stacking at 5:52 AM. MCenter resumed taking beam.

Operators interlocked the MINOS enclosure, which had dropped, at 6 AM.

Operations resumed sending beam to NuMI at 6:28 AM.

At 8:28 AM, Operations halted the NuMI beam due to problems with the helium flow rate. NuMI resumed taking beam, at low intensity, at 8:45 AM.

Booster experts replaced a controls module for EAPS at 11:23 AM.

At 1:19 PM, Operations reported emittance problems in Pbar.

At 3:09 PM, the Booster notcher (B:NOTCH02) tripped off. It reset okay.

At 4:51 PM, NuMI tripped off due to a current sag on a corrector magnet power supply (V108). It resumed normal operations on its own.

At 5:08 PM, Booster suffered a beam problem due to equipment failure. Operators fixed the device and Pbar resumed stacking about twenty minutes later.

At 6:34 PM, the NuMI beam permit tripped off when operators began work on improving Accumulator cooling. NuMI tripped off when operators turned off Pbar beam. It reset okay.

Operations terminated store 4135 at 9:02 PM.

Operations put the TeV into shot setup at 9:15 PM.

Operations established store 4136 at 11:44 PM with an initial luminosity of 102.77E30. Pbar resumed stacking.

MCenter resumed taking beam at 11:48 PM.

At 11:58 PM, the NuMI permit tripped off for the first of three times in two minutes due to Pbar tripping off. Everything reset okay.

Wednesday May 11
The midnight shift began with Operations monitoring store 4136, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

At 12:08 AM, Pbar tripped off and caused NuMI to trip off. It reset okay.

At 12:36 AM, NuMI tripped off due to current on a power supply (E:V108) going low. It reset okay.

At 2:02 AM, BRF16 tripped off for the first of two times in twenty-five minutes. It reset both times.

At 6:27 AM, BRF18 tripped off and would not reset. The indication was a Power Amplifier LCW leak.

The Plan for Wednesday
The plan for today is to continue stacking and stashing.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 5/11/2005   email Fermilab